You are standing on a dff overlooking a lake and decide to throw a tennis ball to your friends in the water below. You throw the tennis bal with a velocity of 17.5 mis at an angle of 31.5 above the horizontal. When the tennis ball leaves your hand it is 12.5 m above the water. How far does the tennis bal travel horizontally before it hits the water? In your calculations, neglect any effects of air resistance. Number
Answer #1

here,

the initial velocity , u = 17.5 m/s

theta = 31.5 degree

h = 12.5 m

let the time taken to hit the water be t

for vertical direction

- h = u * sin(theta) * t - 0.5 * g * t^2

- 12.5 = 17.5 * sin(31.5) * t - 0.5 * 9.81 * t^2

solving for t

t = 2.78 s

the horizontal distance travelled , x = u * cos(theta) * t

x = 17.5 * cos(31.5) * 2.78 m

x = 41.5 m